OpenStudy (anonymous):

(1-9/4)(1-9/25)(1-9/64)......(1-9/400) Does anyone know how to solve this kind of problems?

OpenStudy (anonymous):

First try to find the law of the numbers' change. In the numbers above, we can find that 4, 25, 64 are 2^2, 5^2, 8^2.

OpenStudy (anonymous):

Yes i got that but i can never create an equation to solve the problem

OpenStudy (anonymous):

Give me a minute please.

OpenStudy (anonymous):

All right. We can find that 2+3=5, 5+3=8, so that denominator of next number is 8+3=11, then 14, 17, 20, and 20^2=400.

OpenStudy (anonymous):

yes but is there a way quick way to solve these

OpenStudy (anonymous):

Let me see.

OpenStudy (anonymous):

OK. The denominators can be written as (2+0*3)^2, (2+1*3)^2, (2+2*3)^2, (2+3*3)^2, (2+4*3)^2, (2+5*3)^2, (2+6*3)^2. (2+0*3)^2 =(2+0*3)(2+0*3) =2*2+2(0*3)+2(0*3)+(0*3)*(0*3) =2*2+(4+0*3)(0*3) So there are two 2, and 2*2+0*3 0*3. (2+1*3)^2 =(2+1*3)(2+1*3) =2*2+2(1*3)+2(1*3)+(1*3)(1*3) =2*2+(4+1*3)(1*3) Thus we get this formula. (2+3x)^2=4+(4+3x)(3x)

OpenStudy (anonymous):

@Mathstudent97

OpenStudy (anonymous):

i think it has something to do with 

OpenStudy (anonymous):

Yes. 3 is important.

OpenStudy (anonymous):

In 4, x=0; in 21,x=1; in 64,x=2; then x=3, 4, 5, 6. The formula is (2+3x)^2=4+(4+3x)(3x). (4+3x)(3x)=12x+9x^2 Now first find how many 12 there, then how many 9 there.

OpenStudy (anonymous):

As x=0, 1, 2, 3, 4, 5, 6, so there are (0+1+2+3+4+5+6) 12.

OpenStudy (anonymous):

(0+1+2+3+4+5+6)*12 =21*12 =252

OpenStudy (anonymous):

thanks but i think I"l have to ask my teacher

OpenStudy (anonymous):

I have not completed yet.

OpenStudy (anonymous):

And there are (0^2+1^2+2^2+3^2+4^2+5^2+6^2+7^2) 9. 140*9=1260 And there are 7 numbers, so 7*(1-9)=7*(-8)=-56 the answer is -56/252+1260= -28/756 =-27

OpenStudy (anonymous):

@Mathstudent97